# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 4.2.2. Neck-stretching
Let $(Y,\xi)$ be the contact manifold $M\#\partial W_{0}$ in Theorem 4.1 with the additional assumption that $M$ is Liouville fillable, and $0\neq\beta\in\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$. Then there exist a contact form $\alpha$ for $\xi$, an admissible pair $(H,J)$ of a Hamiltonian and an almost complex structure, a Morse function $f$ on $Y$, a $\mathbb{Q}$-linear combination $\sum a_{i}x_{i}$ of non-constant Hamiltonian orbits of $H$ (of Conley-Zehnder index $2$ when it makes sense, i.e. $c_{1}(M\#\partial W_{0})=0$), and a $\mathbb{Q}$-linear combination $\sum b_{i}p_{i}$ of index $n$ critical points of $f$, such that the following properties are satisfied:
@@
    $\sum b_{i}p_{i}$ represents $0\oplus\beta$ in $H^{n}(M;\mathbb{Q})\oplus H^{n}(\partial W_{0};\mathbb{Q})=H^{n}(Y;\mathbb{Q})$ via Morse cohomology of $f$;
@@
    $\sum a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)=0$, where $y$ is a non-constant Hamiltonian orbit; to be precise, by this we mean that the moduli spaces of the type ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)$ that are of expected dimension at most 00 are cut out transversely, and the sum $\sum a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y)$ for those with expected dimension 00 vanishes;
@@
    ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)=\emptyset$ for all $i$, $y$ and $\Gamma\neq\emptyset$;
@@
    ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{i}},\Gamma)=\emptyset$ for $\Gamma\neq\emptyset$ and $\sum_{i,j}a_{i}b_{j}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{j}},\emptyset)\neq 0$, where $S_{p}$ denotes the stable manifold of $p$ w.r.t. $\nabla f$.
Proof.
For the reader’s convenience and because the two situation require different arguments, we divide the proof in two cases. First, we look at the setting where $W_{0}$ is flexible (of dimension $22n+$ 2) and $c_{1}$ of the contact structure at the contact connected sum vanishes. Here, we use index arguments to understand whether the moduli spaces considered are (generically) empty. Then, we deal with the case of $W_{0}$ subcritical and no assumption on $c_{1}$ of the contact structure on $M$. In this case, the Conley-Zehnder index is (potentially) not well defined, so we argue instead using topological and energy arguments thanks to the specific topological splitting of the form $W_{0}=V\times\mathbb{D}^{2}$ given by [Cie02b].
Case 1: $W_{0}$ is flexible Weinstein (of dimension $22n+$ 2) and $c_{1}(M)=0$.
By [BEE12, MS18], we have that $SH^{*}_{+}(W_{0})\to H^{*}(W_{0})[1]$ is an isomorphism for the standard flexible filling $W_{0}$. Moreover, according to [Laz20a], there exists a contact form $\alpha$ on $\partial W_{0}$ such that all Reeb orbits of period $<D$ are non-degenerate and have Conley-Zehnder indices at least $1$, with the additional property that $SH_{+}^{*,<D}(W_{0})\to H^{*}(W_{0})[1]$ is surjective. In the connected sum, we can assume the contact form on $M$ is large enough so that all Reeb orbits on $M$ have a period $\gg D$. Following [Laz20a, Theorem 3.15] and [Yau04], we can then assume that all the Reeb orbits of period $<D$ on the connected sum $M\#\partial W_{0}$ are those in the connected-summand $\partial W_{0}$, plus the multiple covers $\gamma_{h}^{k}$ of a Reeb orbit $\gamma_{h}$ winding around the core of the $1$-handle, for $k$ smaller than a certain $k_{0}$ depending on $D$ and the “thickness” of the contact handle realising the connected sum. Moreover, the Conley-Zehnder index of $\gamma_{h}^{k}$ is $n+2k-1$.
Let then $F$ be the Liouville filling of $M$. Using the above contact form, we have $H^{n}(W_{0})\to SH^{n,<D}(W_{0})=SH^{n,<D}(F\natural W_{0})$ is zero, where the last equality is given by [Cie02a, Fau20]. As a consequence, we have $SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})$ is surjective onto $0\oplus H^{n}(W_{0})$. Therefore we have $\delta_{\partial}:SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})\to H^{n}(M\#\partial W_{0})$ is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$. Since $\beta\in\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$, there are $x_{i}$ Hamiltonian orbits with Conley-Zehnder index $2$ and a linear combination $\sum a_{i}x_{i}$ representing an element in the positive symplectic cohomology (i.e. in particular closed w.r.t. the differential) so that $\delta_{\partial}(\sum a_{i}x_{i})=\beta$. We moreover choose a Morse function $f$ on $Y$ such that $0\oplus\beta\in H^{n}(M)\oplus H^{n}(\partial W_{0})=H^{n}(Y)$ is represented by $\sum b_{i}p_{i}$, for some index $n$ critical points $p_{i}$ of $f$.
Now we apply neck-stretching to the moduli spaces defining the differential and $\delta_{\partial}$ in order to deduce the desired properties. The expected dimension for ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)$ is
$\mu_{CZ}(x_{i})-\mu_{CZ}(y_{i})-1-\sum_{\gamma\in\Gamma}(\mu_{CZ}(\gamma)+n-2)\leq 2-1-1-\sum_{\gamma\in\Gamma}(n-1)=-(n-1)\,\#\Gamma$
which is negative if $\Gamma\neq\emptyset$; hence (3) follows. Similarly, the expected dimension of ¯$\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{i}},\Gamma)$ is
which is again negative if $\Gamma\neq\emptyset$, hence the first claim of (4) holds. When $y$ has Conley-Zehnder index larger than $1$, (2) follows from dimension counting. When $y$ has Conley-Zehnder index $1$, (2) holds as $\sum a_{i}x_{i}$ is closed for any sufficiently stretched almost complex structure by [Zho21b, Lemma 2.15] and there are no other degenerations in the fully stretched case by (3). For the second claim of (4), we apply neck-stretching to the moduli space contributing to $\delta_{\partial}(\sum a_{i}x_{i})$. Since we can rule out ¯$\overline{{\mathcal{M}}}_{Y,H}(y,S_{p_{i}},\Gamma)$ for any $y$ with Conley-Zehnder index at most $1$ by dimension counting, there cannot be other degeneration in the neck-stretching, and so $\sum_{i,j}a_{i}b_{j}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p_{j}},\emptyset)\neq 0$ corresponds to $\delta_{\partial}(\sum a_{i}x_{i})=0\oplus\beta$, which holds for any sufficiently stretched almost complex structure by [Zho21b, Proposition 2.17].
Case 2: $W_{0}$ is subcritical (and $c_{1}(M)$ is not necessarily zero).
As already pointed out at the beginning of the proof, the lack of the $c_{1}=0$ assumption for the contact structure on $M$ costs us the well-definedness of the Conley-Zehnder indices. We then argue differently, using more the topology of this situation, which allows to derive some crucial energy bounds. More precisely, according to [Cie02b], if $W_{0}$ is subcritical then it splits as a symplectic product as $W_{0}=V\times\mathbb{D}^{2}$, for some Weinstein domain $V$. This product structure and the fact that the considered almost complex structures are compatible with this splitting allows us to get the following lower energy bound for (parts of) Floer cylinders which are entirely contained in the symplectization of the piece $\partial V\times\mathbb{D}^{2}\subset\partial(V\times\mathbb{D}^{2})$, where the chosen Hamiltonian only depends on the symplectization direction $t$: the composition $\pi_{D}\circ u$, for $\pi_{D}\colon\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}\to\mathbb{D}^{2}$ the natural projection, is a holomorphic map to $\mathbb{D}$ and we have
$\int u^{*}\mathrm{d}\lambda\geq\int(\pi_{\mathbb{D}}u)^{*}\mathrm{d}\left(\frac{r^{2}}{2\pi}\mathrm{d}\theta\right)\geq 0\,.$ (*)
The aim is then to use this energy bound (*4) to make up for the loss of the well-definedness of the Conley-Zehnder indices in this case, and still prove the desired conclusion. In order to do so, we need to carefully choose a convenient contact form for the contact structure on $M\#\partial(V\times\mathbb{D}^{2})$.
For this, we start by describing an explicit choice of contact form on $\partial(V\times\mathbb{D}^{2})$ as in [Zho22a, §2.1]. More precisely, one first rounds the corner of
$\left(\,\partial(V\times\mathbb{D}^{2})=V\times S^{1}\cup\partial V\times\mathbb{D}^{2}\,,\;\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta\,\right)\,,$
and the desired contact form $\alpha$ on $\partial(V\times\mathbb{D}^{2})$ is then a perturbation of $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ supported on the mapping torus region $V\times S^{1}$. As explained in detail in [Zho22a, §2.1], this perturbation is achieved via an auxiliary Morse function $f$ on $V$, satisfying the following properties:
  • it is self-indexing, i.e. $f(p)>f(q)$ if and only if $\mathrm{ind}(p)>\mathrm{ind}(q)$ for every pair of critical points $p,q$;
  • $f(p)$ is approximately $1$ for every critical point $p$ of $f$;
  • using $r$ as collar coordinate given by the Liouville vector field near the boundary, $\frac{\partial}{\partial r}(\frac{1}{f})\to\infty$ and $f\to 1$ when approaching the boundary of $V$.
Leaving out the details of the perturbation, for our purposes it is enough to point out that, for an $f$ chosen as above, the perturbed contact form $\alpha$ is given by $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ on $\partial V\times\mathbb{D}^{2}$, and by $\lambda_{V}+\frac{1}{2\pi f}\mathrm{d}\theta$ on $V\times S^{1}=\partial(V\times\mathbb{D}^{2})\,\setminus\,\partial V\times\mathbb{D}^{2}$. In particular, it satisfies the following properties.
@@
    Each critical point $p$ of $f$ corresponds to a simple non-degenerate Reeb orbit $\gamma_{p}$, which is the circle over $p$ in the region $V\times S^{1}\subset\partial(V\times\mathbb{D}^{2})$. In particular, the Reeb orbits $\gamma_{p_{i}}$’s wind around the binding $\partial V\times\{0\}\subset\partial V\times\mathbb{D}^{2}$ exactly once.

### 14 Mar 2026 v5

@@ 4.5. Reeb dynamics and symplectic cohomology of connected sums
Assume $W_{1}$ and $W_{2}$ are $(2n+2)$-dimensional Liouville fillings of $Y_{1}$ and $Y_{2}$ respectively for $n\geq 2$. Let $W_{1}\natural W_{2}$ be the strict Liouville filling of $(Y_{1}\#Y_{2},\alpha_{1}\#_{D}\alpha_{2})$ from Lemma 4.3 (for any $D$ such that Lemma 4.3 applies) such that $c_{1}(W_{1})=c_{1}(W_{2})=0$. Then we have
$SH^{n,<D}(W_{1}\natural W_{2})\simeq SH^{n,<D}(W_{1})\oplus SH^{n,<D}(W_{2})$
via the Viterbo transfer map.
Proof.
We give two explanations of the result, as we believe both are of interest for different reasons: the first uses the specific degree in the statement (i.e. the $SH^{*}$ grading $n$, or equivalently, Conley-Zehnder index $1$); the second uses the general strategy of [Cie02b].
Proof 1: The Conley-Zehnder index of $\gamma_{h,i}^{k}$ is at least $n+1\geq 3$. Hence, all these orbits have cohomological grading [^2] I.e. the degree in $SH^{*}$, namely $n+1-\mu_{CZ}$. at most 0. Therefore $W_{1}\natural W_{2}$ and $W_{1}\sqcup W_{2}$ have the same generators with cohomological grading $n,n+1$. They also share the same generators with cohomological grading n1n-1 if $n\geq 3$ and when $n=2$, $W_{1}\natural W_{2}$ has an extra grading $1$-generator corresponding to the Morse critical point in the $1$-handle of $W_{1}\natural W_{2}$. Using the action filtration as in [Zho23a, §4.2], the Viterbo transfer map from $W_{1}\natural W_{2}$ to $W_{1}\sqcup W_{2}$ induces an isomorphism on positive cochains (i.e. those with generators from Reeb orbits) of cohomological grading n1,n,n+1n-1,n,n+1. Therefore the Viterbo transfer map from $W_{1}\natural W_{2}$ to $W_{1}\sqcup W_{2}$ induces a cochain isomorphism of cohomological grading $n,n+1$ and surjection on n1n-1, which then gives an isomorphism $SH^{n,<D}(W_{1}\natural W_{2})\simeq SH^{n,<D}(W_{1})\oplus SH^{n,<D}(W_{2})$.
Proof 2: Alternatively, the filtered symplectic cohomology can be defined using any Hamiltonian whose slope is the period threshold on the cylindrical end, see [Zho21a, Proposition 2.8]. Thanks to [Cie02a, Lemma 2.5] or [Fau20, §3.3], one can then find a Hamiltonian on $W_{1}\natural W_{2}$ defining $SH^{*,<D}(W_{1}\natural W_{2})$, such that the generators are those of $W_{1}\sqcup W_{2}$ defining filtered symplectic cohomology of $W_{1}\sqcup W_{2}$ along with one extra generator – a critical point – in the $1$-handle with Conley-Zehnder index $\gg n$ [^3] The Conley-Zehnder index of such a generator is actually $n+2k_{0}n$, where $k_{0}$ is as in Lemma 4.3. This index can be understood as the index of the survivor in cohomology of the complex of check and hat orbits from (2) of Lemma 4.3, which is the hat version of the largest index in (2) of Lemma 4.3. The Hamiltonian is not $C^{2}$-small near the critical points. In the context of 0 handles, this is [Sei08, (3f)]. The case of $1$-handles is precisely the product of a codimension 22 0-handle and $\mathbb{C}$ with a hyperbolic point. . That is the unique critical point in the handle with a large positive Hessian in a codimension $2$ symplectic subspace and a hyperbolic Hessian in the normal direction, see [Cie02a, Lemma 2.5 with suitable $A_{i},B_{i}$]. By [Cie02a, Theorem 1.11, before taking the limit w.r.t. slopes], we get isomorphism
$SH^{n,<D}(W_{1}\natural W_{2})\simeq SH^{n,<D}(W_{1})\oplus SH^{n,<D}(W_{2}).$
The second approach works for any degree, not just $n$, see [CO18, Proposition 9.19], where the statement is for full symplectic cohomology while the proof factors through the filtered version. ∎
4.6. Moduli spaces for the connecting map in Theorem 4.1
We are now ready to state our main technical result, wherein the neck-stretching is used to analyze precisely what curves appear in the computation of the connecting map $\delta_{\partial}\colon SH_{+}^{*,<D}(W,\Lambda)\to H^{*}(Y,\Lambda)[1]$ under the hypotheses of Theorem 4.1. Here, $\Lambda$ is the Novikov field, which appears due to the fact that we consider general strong fillings. Moreover, despite the fact that we will consider strong fillings of the connected sum, we will use a cobordism trick to reduce to the case that $M$ is Liouville (in fact, Weinstein) fillable.
Proposition 4.5.
Let $(Y,\xi)$ denote be the contact connected sum $M\#\partial W_{0}$ given in the statement of Theorem 4.1 with the additional assumption that $M$ is Liouville fillable and the Liouville filling has trivial first Chern class in the case (a) of Theorem 4.1, and $\beta$ a non-trivial class in the image of $H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q})$. Then there are
  • a contact form $\alpha$ for $\xi$,
  • an admissible pair $(H,J)$ of a Hamiltonian and an almost complex structure on the completion $\widehat{Y}$ as in §4.4,
  • a Morse function $g$ on $Y$,
  • a $\mathbb{Q}$-linear combination $\sum_{i}a_{i}x_{i}$ of non-constant Hamiltonian orbits of $H$, which are of Conley-Zehnder index $2$ in the case that $c_{1}(M\#\partial W_{0})=0$,
so that the following properties hold:
@@
    $\sum_{i}a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\emptyset)=0$, for every non-constant Hamiltonian orbit $y$; to be precise, the moduli spaces of the type $\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\emptyset)$ that are of expected dimension at most 0 are cut out transversely, and the sum $\sum a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\emptyset)$ for those with expected dimension 0 vanishes;
@@
    $\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)=\emptyset$ for all $i$, $y$ and $\Gamma\neq\emptyset$ multiset of $\alpha$-Reeb orbits;
@@
    For any index $n$ critical point $p$ of $g$, we have $\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p},\Gamma)=\emptyset$ for $\Gamma\neq\emptyset$. Moreover, $\sum_{i}\sum_{p}a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p},\emptyset)\,p$ represents $0\oplus\beta$ in $H^{n}(M;\mathbb{Q})\oplus H^{n}(\partial W_{0};\mathbb{Q})=H^{n}(Y;\mathbb{Q})$ in the Morse cochain complex of $g$, where the inner sum is over all index $n$ critical points $p$ of $g$ and $S_{p}$ denotes the stable manifold of $p$ with respect to the gradient $\nabla g$.
The above result is what guarantees that cohomology classes in the image of $H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q})$ are “detected” via the connecting map $\delta_{\partial}$ and is independent of the strong filling. Indeed, the significance of the properties above is the following: (1) combined with (2) will tell us that $\sum_{i}a_{i}x_{i}$ is closed with respect to the Floer differential and defines an element in the positive symplectic cohomology with respect to any symplectic filling; (3) will imply that the degree-$n$ term in $\delta_{\partial}(\sum a_{i}x_{i})$ is exactly $0\oplus\beta$.
We postpone the proof of Proposition 4.5, and use it first to prove Theorem 4.1.
Proof of Theorem 4.1.
As in the statement, assume $W$ is a strong filling of $Y$. By the universal coefficient theorem, it is equivalent to proving the dual statement that $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$ is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$. By [Laz20b, Theorem 1.7], there is a Weinstein cobordism from $M$ to $M^{\prime}$, such that $M^{\prime}$ is Weinstein fillable. In the case where $\dim M=5$, according to [BCS15, Proposition 3.4], we can moreover assume that the Weinstein cobordism has vanishing first Chern class if $M$ does. In particular, $M^{\prime}$ has a Weinstein filling with vanishing first Chern class. Then, we have a Weinstein cobordism from $Y$ to $Y^{\prime}:=M^{\prime}\#\partial W_{0}$, yielding a strong filling $W^{\prime}$ of $Y^{\prime}$. Observe that the desired claim follows from the corresponding statement for $Y^{\prime}=\partial W^{\prime}$, since the surgeries do not affect the $\partial W_{0}$-factor. In other words, it suffices to prove the conclusion of Theorem 4.1 under the additional assumption that $M$ is Liouville fillable.
Let $\beta\in\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$; we need to prove that $0\oplus\beta$ is in the image of $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$. As we can assume that $M$ has a Liouville filling $F$, we can apply Proposition 4.5 and we let $\sum a_{i}x_{i}$ be the resulting $\mathbb{Q}$-linear combination of non-constant periodic orbits of the Hamiltonian. Viewing $H$ as a Hamiltonian on $\widehat{W}$ by extending it as zero over the filling, we can then consider the sum $\sum a_{i}q^{\mathcal{A}_{H}(x_{i})}x_{i}$ as an element of the Floer complex of $H$ (considered as a Hamiltonian on $\widehat{W}$), with coefficients in the Novikov field [^4] Symplectic cohomology of strong fillings requires Novikov coefficients, and different orders of applying direct limit and completion in the construction yield different theories, c.f. [Ven18]. As we only use symplectic cohomology of a finite slope, such subtleties do not arise. $\Lambda$, where $q$ is the formal variable in $\Lambda$ and the symplectic action $\mathcal{A}_{H}(x_{i})$ is computed using the Liouville form in the cylindrical end containing $x_{i}$.
We now claim that, for a sufficiently stretched almost complex structure, the chain $\sum a_{i}q^{\mathcal{A}_{H}(x_{i})}x_{i}$ is in fact closed w.r.t. the Floer differential so that the sum represents a class in $SH_{+}^{*}(W;\Lambda)$. Indeed, using neck-stretching along the contact boundary, by (1) and (2) of Proposition 4.5, one can arrange that all the Floer cylinders used in the differential are contained in the positive end of the symplectization, whose energy is given by the difference of the respective symplectic actions $\mathcal{A}_{H}$, which is well-defined on the positive end of the completion. Therefore, using (1) of Proposition 4.5, we have that
$\begin{split}\delta\left(\sum_{i}a_{i}q^{\mathcal{A}_{H}(x_{i})}x_{i}\right)=&\sum_{y}\sum_{i}a_{i}\#\overline{\mathcal{M}}_{Y,H}(x_{i},y,\emptyset)q^{\mathcal{A}_{H}(x_{i})+\mathcal{A}_{H}(y)-\mathcal{A}_{H}(x_{i})}y\\ =&\sum_{y}\underbrace{\left[\sum_{i}a_{i}\#\overline{\mathcal{M}}_{Y,H}(x_{i},y,\emptyset)\right]}_{=\,0}q^{\mathcal{A}_{H}(y)}y=0.\end{split}$
Similarly, (3) of Proposition 4.5 implies that $\sum a_{i}q^{\mathcal{A}_{H}(x_{i})}x_{i}$ is mapped to $0\oplus\beta$ under the composition
$\Pi_{n}\circ\delta_{\partial}\colon SH^{*}_{+}(W;\Lambda)\to H^{*+1}(W;\Lambda)\to H^{*+1}(Y;\Lambda)\to H^{n}(Y;\Lambda)\,$,
where $\Pi_{n}\colon H^{*+1}(Y;\Lambda)\to H^{n}(Y;\Lambda)$ denotes the natural projection and $0\oplus\beta\in H^{*}(Y;\mathbb{Q})$ is considered as an element in $H^{*}(Y;\Lambda)\cong H^{*}(Y;\mathbb{Q})\otimes_{\mathbb{Q}}\Lambda$. Hence, $0\oplus\beta$ also lies in the image of $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$, as desired. ∎
Proof of Proposition 4.5 when $c_{1}(Y)=0$
Roughly speaking, the proof is structured as follows: We first describe how to find the contact form $\alpha$ on $Y$, an admissible pair $(H,J)$, and the $\mathbb{Q}$-linear combination $\sum_{i}a_{i}x_{i}$ as in the statement. The auxiliary Morse function is arbitrary. Then, we proceed to prove all the properties of these objects as claimed in the statement. For this, the strategy of stretching the neck will be used repeatedly. We now give the details.
Proof of Proposition 4.5 for $c_{1}(Y)=0$.
By the surgery formulae in [Cie02a, Theorem 1.11] and [BEE12, Theorem 5.6], or alternatively [MS18, Theorem 3.2], the symplectic cohomology $SH^{*}(W_{0})$ is trivial so that the connecting map $SH^{*}_{+}(W_{0})\to H^{*}(W_{0})[1]$ is an isomorphism for the flexible filling $W_{0}$. Moreover, according to [Laz20a, Theorem 1.9], there is a sequence of contact forms $\alpha_{1}>\alpha_{2}>\ldots$ and positive numbers $D_{i}\to\infty$, such that all contractible Reeb orbits $\alpha_{i}$ with period smaller than $D_{i}$ are non-degenerate and have Conley-Zehnder index at least 11555[Laz20a, Theorem 1.9] states that the SFT degrees are at least $1$, but the proof actually shows that Conley-Zehnder indices are at least $1$ by [Laz20a, Propositions 4.3, 4.4], see [Zho23c, Example 2.6 (1)].. As the direct limit of symplectic cohomology of $W_{0}$ with boundary contact form $\alpha_{i}$ and period upper threshold $D_{i}$, through the Viterbo transfer maps and continuation maps, is the symplectic cohomology of $W_{0}$, see e.g. [Laz20a, Proof of Proposition 3.8], we may assume there exists a contact form $\alpha_{0}$ on $\partial W_{0}$ and a positive real number $D$ such that all Reeb orbits of period <D<D are non-degenerate and have Conley-Zehnder indices at least $1$, with the additional property that $SH_{+}^{*,<D}(W_{0})\to H^{*}(W_{0})[1]$ is surjective, or equivalently $H^{*}(W_{0})\to SH^{*,<D}(W_{0})$ is zero.
We now consider the connected sum of the strict contact manifolds $(\partial W_{0},\alpha_{0})$ and $(M,K\alpha_{M})$, for some $K\gg 0$, and contact form $\alpha_{M}$ on $M$. According to Lemma 4.3, as long as $K\gg D$, we can assume that all the Reeb orbits of period <D<D on the connected sum $(M\#\partial W_{0},\alpha)$ are those in the connected-summand $\partial W_{0}$, plus some of the multiple covers $\gamma_{h,i}^{k}$ of simple Reeb orbit $\gamma_{h,i}$ on the co-core of the $1$-handle. Let now $F$ be the Liouville filling of $M$ with trivial first Chern class. Using the explicit contact form on the boundary $M\#\partial W_{0}$ of the boundary connect sum $F\natural W_{0}$ as described above, Proposition 4.4 implies that
$SH^{n,<D}(F\natural W_{0})=SH^{n,<D}(F)\oplus SH^{n,<D}(W_{0})$
via the Viterbo transfer map. As a consequence, we know that $H^{n}(W_{0})\to SH^{n,<D}(W_{0})\to SH^{n,<D}(F\natural W_{0})$ is zero.
Now, the long exact sequence for filtered positive symplectic cohomology implies that $SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})$ is surjective onto $0\oplus H^{n}(W_{0})$. Therefore, we have that
$\delta_{\partial}:SH^{n-1,<D}_{+}(F\natural W_{0})\to H^{n}(F\natural W_{0})\to H^{n}(M\#\partial W_{0})$
is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$. Since $\beta\in\operatorname{image}(H^{n}(W_{0})\to H^{n}(\partial W_{0}))$, this then implies that there are $x_{i}$ Hamiltonian orbits with Conley-Zehnder index $2$ and a linear combination $\sum a_{i}x_{i}$ representing an element in the positive symplectic cohomology (i.e. in particular closed w.r.t. the differential) so that $\delta_{\partial}(\sum a_{i}x_{i})=\beta$.
Now we apply neck-stretching to the moduli spaces defining the differential and $\delta_{\partial}$ in order to deduce the desired properties. The expected dimension for $\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\Gamma)$ is
$\mu_{CZ}(x_{i})-\mu_{CZ}(y)-1-\sum_{\gamma\in\Gamma}(\mu_{CZ}(\gamma)+n-2)\leq 2-1-1-\sum_{\gamma\in\Gamma}(n-1)=-(n-1)\,\#\Gamma$
which is negative if $\Gamma\neq\emptyset$; hence (2) follows. Similarly, the expected dimension of $\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p},\Gamma)$ is
which is again negative if $\Gamma\neq\emptyset$, hence the first claim of (3) holds.
When $y$ has Conley-Zehnder index larger than $1$, (1) holds since the expected dimension in this case is negative. When $y$ has Conley-Zehnder index $1$, by (2), for a sufficiently stretched almost complex structure, the Floer cylinders that are counted when defining $SH^{*}_{+}(F\natural W_{0})$ are completely contained in the positive end of $\widehat{F\natural W_{0}}$, i.e. the corresponding moduli space can be identified with $\overline{{\mathcal{M}}}_{Y,H}(x_{i},y,\emptyset)$. Since the above properties on filtered symplectic cohomology do not depend on the choice of almost complex structure, we may assume $\sum a_{i}x_{i}$ is closed for a sufficiently stretched almost complex structure. This implies (1).
For the second claim of (3), we apply neck-stretching to the moduli space contributing to $\delta_{\partial}(\sum a_{i}x_{i})$. Since we have $\overline{{\mathcal{M}}}_{Y,H}(y,S_{p},\Gamma)=\emptyset$ for any $y$ with Conley-Zehnder index at most $1$ by dimension counting, there cannot be other degeneration in the neck-stretching, and so for any sufficiently stretched almost complex structure the curve contributing to $\delta_{\delta}(\sum a_{i}x_{i})$ are completely contained in the positive end of $\widehat{F\natural W_{0}}$, which can be identified with ${\mathcal{M}}_{Y,H}(x_{i},S_{p},\emptyset)$. Therefore we have
$\sum_{i}\sum_{p}a_{i}\,\#\overline{{\mathcal{M}}}_{Y,H}(x_{i},S_{p},\emptyset)p=0\oplus\beta\in H^{n}(M\#\partial W_{0}).$
4.7. The subcritical case
As already pointed out above, the fact that $c_{1}(Y)$ need not vanish means that the Conley-Zehnder indices are not well-defined. In this case, we make up for this loss with an energy bound (in (**3) below), which is derived leveraging the subcriticality of $W_{0}$ in a substantial way.
Geometric set-up
We start by describing an explicit choice of contact forms on $\partial(V\times\mathbb{D}^{2})$ as in [Zho23b, §2.1]. More precisely, one first rounds the corner of
$\left(\,\partial(V\times\mathbb{D}^{2})=V\times\mathbb{S}^{1}\cup_{\partial V\times\mathbb{S}^{1}}\partial V\times\mathbb{D}^{2}\,,\;\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta\,\right)\,$, (9)
and the desired contact form $\alpha$ on $\partial(V\times\mathbb{D}^{2})$ is then a perturbation of $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ supported on the mapping torus region $V\times S^{1}$. As explained in detail in [Zho23b, §2.1], this perturbation is achieved via an auxiliary Morse function $f$ on $V$, satisfying the following properties:
  • it is “self-indexing”, by which we mean that $f(p)>f(q)$ if and only if $\mathrm{ind}(p)>\mathrm{ind}(q)$ for every pair of critical points $p,q$ (in particular, all critical points of the same index have the same image under $f$);
  • $f(p)$ is approximately $1$ for every critical point $p$ of $f$;
  • $f\to 1$ when approaching the boundary of $V$, whose precise meaning will be clear from the discussion below.
We use such $f$ to smooth the corners of the hypersurface $\partial(V\times\mathbb{D}^{2})$ inside the completion $(\widehat{V}\times\mathbb{C},\widehat{\lambda}_{V}+\frac{r^{2}}{2\pi}{\mathrm{d}}\theta)$ by bumping up the $V\times\mathbb{S}^{1}$ using the graph of $r^{2}=1/f$. It is clear that we can choose $f$ such that ${\mathrm{d}}f$ is $C^{1}$-small outside a collar neighborhood of $\partial V$ in $V$ and $f$ only depends on the Liouville coordinate near $\partial V$, whose derivatives diverge when approaching the boundary, to make the perturbed hypersurface smooth and of contact type. For an $f$ chosen as above, the perturbed contact form $\alpha$ is given by $\lambda_{V}+\frac{r^{2}}{2\pi}\mathrm{d}\theta$ on $\partial V\times\mathbb{D}^{2}$, and by $\lambda_{V}+\frac{1}{2\pi f}\mathrm{d}\theta$ on $V\times\mathbb{S}^{1}=\partial(V\times\mathbb{D}^{2})\,\setminus\,\partial V\times\mathbb{D}^{2}$. These are explained in detail in [Zho23b, §2.1]. Moreover, it satisfies the following properties.
@@ 4.7. The subcritical case
    Each critical point $p$ of $f$ corresponds to a simple non-degenerate Reeb orbit $\gamma_{p}$, which is the circle over $p$ in the region $V\times\mathbb{S}^{1}\subset\partial(V\times\mathbb{D}^{2})$. In particular, the Reeb orbits $\gamma_{p_{i}}$’s wind around the binding $\partial V\times\{0\}\subset\partial V\times\mathbb{D}^{2}$ exactly once.
@@
    The period of $\gamma_{p}$ is $1/f(p)$ (which is approximately $1$ by our choice of $f$), and hence $\gamma_{p}$ has longer period than $\gamma_{q}$ if and only if $f(p)<f(q)$.